Sunlight at the radius of Earth’s orbit away from the sun contains 1370W atts/m2 (this is called the solar constant and is a very important constant for climate science). i.e. Every second, 1370J of sunlight fly through a square of size 1m2. The Earth has an albedo of 0.3. This means that 30% of the energy gets re- flected back into space. For symplicity, we can assume that the a square meter of the Earth facing the sun absorbs 1000J of energy every second. How much energy does Earth absosb every second?
Answer #1

we can assume that the a square meter of the Earth facing the sun absorbs 1000J of energy every second.

Energy absorbed by earth every second = 1000*4*pi*(radius of earth)^2 =  1000*4*pi* (6.37*10^6) ^2= 5.1 * 10^17 J